# LoraHub: Efficient Cross-Task Generalization via Dynamic LoRA Composition

### v2

@@ 2 Problem Statement
Large Language Models We assume that a large language model is based on Transformer architecture (Vaswani et al., 2017) and has been pre-trained on a large-scale text corpus. The model architecture can be either encoder-decoder (Raffel et al., 2020) or decoder-only (Brown et al., 2020). Also, could also have been fine-tuned with a large set of instruction-following datasets such as Flan Colleciton (Longpre et al., 2023) and PromptSource (Bach et al., 2022).
Cross-Task Generalization In real-world situations, users often desire an LLM to perform novel tasks that it has not encountered before — an ability widely known as cross-task generalization. Generally, cross-task generalization falls into two categories: zero-shot learning (Mishra et al., 2022; Sanh et al., 2022; Chung et al., 2022; OpenAI, 2022; Lin et al., 2022), which necessitates no labeled examples of the new task, and few-shot learning (Ye et al., 2021; Min et al., 2022) which demands a handful of labeled examples. Assume we have distinct upstream tasks that the LLM has been trained on, denoted as . Our paper primarily focuses on the latter category, where for an unseen target task , users can only provide a limited set of labeled examples, . Our aim is to modify the model to adapt it to task using only . An intuitive method would be to fine-tune the weights of based on , yielding an updated model with enhanced performance on . However, this approach is inefficient, time-consuming, and unstable when is small.
LoRA Tuning LoRA (Hu et al., 2022), a parameter-efficient fine-tuning method, facilitates the adaptation of LLMs using lightweight modules, eliminating the need for fine-tuning the entire weights. LoRA tuning involves keeping the original model weights frozen while introducing trainable low-rank decomposition matrices as adapter modules into each layer of the model. Compared to the base LLM, this module possesses significantly fewer trainable parameters, paving the way for rapid adaptation using minimal examples. As such, LoRA tuning presents a resource-efficient technique to quickly adapt LLMs for new tasks with restricted training data. However, traditional LoRA methods primarily concentrate on training and testing within the same tasks (Gema et al., 2023), rather than venturing into few-shot cross-task generalization.
@@ 4.1 Experimental setup
Our methodology requires a compendium of LoRA modules trained on preceding tasks. For parity with FLAN, we adopt the tasks utilized to instruct FLAN-T5, thereby incorporating nearly distinct tasks and their corresponding instructions 111We released used the dataset at huggingface.co/datasets/lorahub/flanv2.. Following this, we trained several LoRA modules as potential candidates. During each experimental sequence, we randomly select LoRA modules from them as the candidate for our LoraHub learning.
@@
| Task | Zero | ICL | IA3 | LoRA | FFT | LoraHub | ||
@@ 5 Experimental Analysis
Can LoraHub work well on non-instruction-tuning models?
In previous investigations, we primarily focused on models with zero-shot capabilities that were trained with instruction tuning. However, for models like T5 without zero-shot abilities, where training has a larger effect on parameters, it was unclear if LoraHub could still effectively manage and improve them. Our experiments show that although these models perform worse than FLAN-T5, LoraHub learning can still enable them to effectively generlize to unseen tasks. See Appendix B for more details.
The parameter rank plays a crucial role in the LoRA framework, directly influencing the number of trainable parameters utilized during LoRA tuning. This prompts an intriguing question: does the variation in rank values influence the outcomes observed within the LoraHub learning? Our analysis indicates that, for FLAN-T5, the choice of rank has minimal impact. However, for T5, it still exerts some influence. Empirical findings reveal that, in comparison to rank values of or , a rank value of consistently demonstrates superior performance across different runs, both in terms of average and optimal values. Additional results are available in Appendix B.
Does more LoRA modules lead to better results?
In our main experiments, we randomly selected LoRA modules for LoraHub learning. Therefore, we conducted experiments to investigate the effect of using different numbers of LoRA modules. The results demonstrate that as we increased the number of LoRA modules, the variance in performance increased. However, the maximum achievable performance also improved. More analysis on the variance and the detailed results can be found in Appendix G.
| LoRA Retrieval | LoraHub | LoraHub |
| 31.7 | 34.7 | 41.2 |
6 Related work
Model Merging
The second category most closely aligns with our research, stemming from a shared motivation of module composition. Various scholars have made advances in this line of research: Kingetsu et al. (2021) decomposes and recomposes modules on the basis of their functionality; Ilharco et al. (2023) proposes modulating model behavior using task vectors; Lv et al. (2023) amalgamates parameter-efficient modules weighted according to task similarity; Zhang et al. (2023) crafts modules by employing specific arithmetic operations; Sun et al. (2023) improves few-shot performance of unseen tasks by multi-task pre-training of prompts; Chronopoulou et al. (2023) averages adapter weights intended for transfer; Ponti et al. (2023) focuses on jointly learning adapters and a routing function that allocates skills to each task; and Muqeeth et al. (2023) concentrates on amalgamating experts in mixture of experts models; However, these methods generally necessitate multi-task training or human prior on module selection for the downstream task. In contrast, our method does not impose any special training requirements and simply employs vanilla LoRA tuning. Additionally, the module selection for downstream tasks is entirely data-driven without human prior knowledge. This design gives the advantage of easily adding new LoRA modules for reuse, allowing our method to flexibly scale up the number of LoRA module candidates in the future.
Mixture of Experts
Cross-Task generalization
7 Conclusion
8 Limitations & future work
Pre-Filtering of LoRA Module Candidates
While our method is successful in identifying and weighting relevant aspects from seen tasks to enhance unseen task performance, relying entirely on the model to perform this search can lead to increased computational demands and potentially unstable results. Incorporating a pre-filtering step to select only pertinent LoRA modules could expedite and refine performance. Identifying an effective selection strategy warrants further study.
Method Applicability to Decoder-Only Models
All experiments for this study were executed using the encoder-decoder architecture. We aspire to extrapolate this method to decoder-only models such as GPT (Brown et al., 2020), aiming to determine its applicability in such contexts.
Exploring Superior Optimization Methods
The use of a genetic algorithm for optimization in this study raises the question of whether better optimization approaches exist that could provide superior gradient-free optimization with limited examples. Although the current method has shown adequate performance, there is still room for improvement.
Appendix A Result of Best Results
@@
| Task | ICL | IA3 | LoRA | FFT | LoraHub | ||
@@
Appendix B Result of non-instrcution-tuned models
| Task Rank | 4 | 4 | 16 | 16 | 64 | 64 | ||
@@ Appendix B Result of non-instrcution-tuned models
Appendix C Result of larger model
| Task | Zero | LoraHub | LoraHub | ||
@@ Appendix C Result of larger model
Appendix D Improving the Robustness of LoraHub
@@
| Task | LoraHub | LoraHub | ||
@@
Appendix E Performance on General Important Task
@@
Appendix F Implementation details
@@
Appendix G Influence of Number of LoRA modules
@@
Appendix H The Impact of Threshold
@@
| Task | LoraHub with threshold | LoraHub without threshold | ||

### v3

@@ 2 Problem Statement
Large Language Models
We assume that a large language model is based on Transformer architecture (Vaswani et al., 2017) and has been pre-trained on a large-scale text corpus. The model architecture can be either encoder-decoder (Raffel et al., 2020) or decoder-only (Brown et al., 2020). Also, could also have been fine-tuned with a large set of instruction-following datasets such as Flan Colleciton (Longpre et al., 2023) and PromptSource (Bach et al., 2022).
Cross-Task Generalization
In real-world situations, users often desire an LLM to perform novel tasks that it has not encountered before — an ability widely known as cross-task generalization. Generally, cross-task generalization falls into two categories: zero-shot learning (Mishra et al., 2022; Sanh et al., 2022; Chung et al., 2022; OpenAI, 2022; Lin et al., 2022), which necessitates no labeled examples of the new task, and few-shot learning (Ye et al., 2021; Min et al., 2022) which demands a handful of labeled examples. Assume we have distinct upstream tasks that the LLM has been trained on, denoted as . Our paper primarily focuses on the latter category, where for an unseen target task , users can only provide a limited set of labeled examples, . Our aim is to modify the model to adapt it to task using only . An intuitive method would be to fine-tune the weights of based on , yielding an updated model with enhanced performance on . However, this approach is inefficient, time-consuming, and unstable when is small.
LoRA Tuning
LoRA is a parameter-efficient fine-tuning method (Hu et al., 2022), facilitates the adaptation of LLMs using lightweight modules, eliminating the need for fine-tuning the entire weights. LoRA tuning involves keeping the original model weights frozen while introducing trainable low-rank decomposition matrices as adapter modules into each layer of the model. Compared to the base LLM, this module possesses significantly fewer trainable parameters, paving the way for rapid adaptation using minimal examples. As such, LoRA tuning presents a resource-efficient technique to quickly adapt LLMs for new tasks with restricted training data. However, traditional LoRA methods primarily concentrate on training and testing within the same tasks (Gema et al., 2023), rather than venturing into few-shot cross-task generalization.
@@ 4.1 Experimental setup
Our methodology requires a compendium of LoRA modules trained on preceding tasks. For parity with FLAN, we adopt the tasks utilized to instruct FLAN-T5, thereby incorporating nearly distinct tasks and their corresponding instructions. Following this, we trained several LoRA modules as potential candidates. During each experimental sequence, we randomly select LoRA modules from them as the candidate for our LoraHub learning.
@@
| Task | Zero | ICLavg | IA3avg | LoRAavg | FFTavg | LoraHubavg | ||
@@ 5 Experimental Analysis
| LoRA Retrieval | LoraHub avg | LoraHub best |
| 31.7 | 34.7 | 41.2 |
How effective is the gradient-free optimization method?
Can LoraHub work well on non-instruction-tuning models?
In previous investigations, we primarily focused on models with zero-shot capabilities that were trained with instruction tuning. However, for models like T5 without zero-shot abilities, where training has a larger effect on parameters, it was unclear if LoraHub could still effectively manage and improve them. Our experiments show that although these models perform worse than FLAN-T5, LoraHub learning can still enable them to effectively generlize to unseen tasks. See Appendix C for more details.
The parameter rank plays a crucial role in the LoRA framework, directly influencing the number of trainable parameters utilized during LoRA tuning. This prompts an intriguing question: does the variation in rank values influence the outcomes observed within the LoraHub learning? Our analysis indicates that, for FLAN-T5, the choice of rank has minimal impact. However, for T5, it still exerts some influence. Empirical findings reveal that, in comparison to rank values of or , a rank value of consistently demonstrates superior performance across different runs, both in terms of average and optimal values. Additional results are available in Appendix C.
Does more LoRA modules lead to better results?
In our main experiments, we randomly selected LoRA modules for LoraHub learning. Therefore, we conducted experiments to investigate the effect of using different numbers of LoRA modules. The results demonstrate that as we increased the number of LoRA modules, the variance in performance increased. However, the maximum achievable performance also improved. More analysis on the variance and the detailed results can be found in Appendix H.
How much computational resource can be saved? We follow to the memory test settings from the LoRA-FA (Zhang et al., 2023b) study for an accurate benchmark. In this context, full fine-tuning required about 40GB of memory, whereas LoRA fine-tuning used around 34GB. Remarkably, LoraHub only utilized about 5GB of memory, illustrating its efficiency due to the inference-only mode, which eliminates the need for storing gradients and optimization states.
6 Related work
Model Merging
The second category most closely aligns with our research, stemming from a shared motivation of module composition. Various scholars have made advances in this line of research: Kingetsu et al. (2021) decomposes and recomposes modules on the basis of their functionality; Ilharco et al. (2023) proposes modulating model behavior using task vectors; Lv et al. (2023) amalgamates parameter-efficient modules weighted according to task similarity; Zhang et al. (2023a) crafts modules by employing specific arithmetic operations; Sun et al. (2023) improves few-shot performance of unseen tasks by multi-task pre-training of prompts; Chronopoulou et al. (2023) averages adapter weights intended for transfer; Ponti et al. (2023) focuses on jointly learning adapters and a routing function that allocates skills to each task; and Muqeeth et al. (2023) concentrates on amalgamating experts in mixture of experts models; However, these methods generally necessitate multi-task training or human prior on module selection for the downstream task. In contrast, our method does not impose any special training requirements and simply employs vanilla LoRA tuning. Additionally, the module selection for downstream tasks is entirely data-driven without human prior knowledge. This design gives the advantage of easily adding new LoRA modules for reuse, allowing our method to flexibly scale up the number of LoRA module candidates in the future.
Mixture of Experts
Cross-Task generalization
7 Conclusion
Reproducibility Statement
The authors have made great efforts to ensure the reproducibility of the empirical results reported in this paper. Firstly, the experiment settings, evaluation metrics, and datasets were described in detail in Section 4.1. Secondly, the codes and script for reproduce the result will be opensource after accepted. Second, the source code implementing the proposed method and experiments will be made publicly available at upon acceptance of the paper. Third, pre-trained LoRA modules from this work along with their configuration files and weights will be shared. These allow reproduction without retraining the LoRA modules, enabling quick testing and verification.
Appendix A More Analysis
@@
Appendix B Result of Best Results
@@
| Task | ICLbest | IA3best | LoRAbest | FFTbest | LoraHubbest | ||
@@
Appendix C Result of non-instrcution-tuned models
| Task Rank | 4avg | 4best | 16avg | 16best | 64avg | 64best | ||
@@ Appendix C Result of non-instrcution-tuned models
Appendix D Result of larger model
| Task | Zero | LoraHub avg | LoraHub best | ||
@@ Appendix D Result of larger model
Appendix E Improving the Robustness of LoraHub
@@
| Task | LoraHubavg | LoraHubfilter | ||
@@
Appendix F Performance on General Important Task
@@
Appendix G Implementation details
@@
Appendix H Influence of Number of LoRA modules
@@
Appendix I The Impact of Threshold
@@
| Task | LoraHubavg with threshold | LoraHubavg without threshold | ||